EXN Status Page
All Systems Operational
Client Portal ? Operational
Shared Hosting ? Operational
DC Connect ? Operational
Wave Connect ? Operational
IP Transit Services ? Operational
Leased Lines Operational
Virtual Data Centre (vDC) - Global Switch 2 ? Operational
Virtual Data Centre (vDC) - Virtus Enfield ? Operational
Datacentres Operational
DRT Dublin Operational
Equinx LD7 Operational
Equinix LD8 Operational
Global Switch 2 Operational
Iron Mountain AMS1 Operational
London Hosting Centre (LHC) Operational
Virtus LON1 Enfield Operational
Volta GSS Operational
Telehouse North Operational
Telehouse East Operational
Virtus LON4 Slough Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Past Incidents
Feb 27, 2024

No incidents reported today.

Feb 26, 2024

No incidents reported.

Feb 25, 2024

No incidents reported.

Feb 24, 2024
Resolved - This incident has been resolved.
Feb 24, 18:04 GMT
Update - The fix was implemented at 14:00 and the link has been stable since. If you are still experiencing issues, please contact EXN support through the usual channels.
Feb 24, 18:04 GMT
Identified - The issue has been identified and a fix is being implemented.
Feb 24, 13:13 GMT
Update - EXN and 3rd party engineers continue to work on this incident. We do not yet have an estimated time of repair.
Feb 24, 09:05 GMT
Update - OTDR results indicate no issues on the dark fibre. We suspect a cross connect fault and have raised this with Equinix
Feb 23, 23:06 GMT
Update - Engineer has arrived at Equinix LD8 and will shortly begin the OTDR test to determine the location of the fault
Feb 23, 18:41 GMT
Update - Our fibre provider has provided an ETA of 18:40 for the engineer to arrive at site
Feb 23, 17:51 GMT
Update - We are waiting for our fibre provider to dispatch an engineer to conduct intrusive testing on this circuit
Feb 23, 15:13 GMT
Investigating - At 13:49 today we were alerted to a fibre flap between Equinix LD8 and Global Switch (2) East in the Docklands area. Since the flap we have seen a degradation of around 10dBm. We are raising this with our dark fibre provider.
Feb 23, 14:07 GMT
Feb 23, 2024
Feb 22, 2024

No incidents reported.

Feb 21, 2024

No incidents reported.

Feb 20, 2024

No incidents reported.

Feb 19, 2024

No incidents reported.

Feb 18, 2024
Completed - The scheduled maintenance has been completed.
Feb 18, 06:30 GMT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Feb 18, 00:30 GMT
Scheduled - Following up from a temporary repair completed to the 24V control systems on 2024-02-05, Verne Global will be applying a permanent fix during this maitenance window.

The new batteries and chargers for the SCADA system are currently fed from temporary power supplies which will be moved to permanent supplies.

The works will be carried out on HV-A and HV-B. Only one system will be worked on at a time and no impact is expected.

Should you have any questions regarding this maintenance, please do contact us through the usual channels.

Feb 16, 09:50 GMT
Feb 17, 2024

No incidents reported.

Feb 16, 2024

No incidents reported.

Feb 15, 2024

No incidents reported.

Feb 14, 2024

No incidents reported.

Feb 13, 2024

No incidents reported.